40ml of hydrocarbon combusted in 260 ml of oxygen obtained 100ml of co2 what is formula of hydrocarbon

Grade:12th pass

1 Answers

Naveen Kumar
9 years ago
The balenced equation can be given as:
CnHm...+(n+m/4)O2............>nCO2.....+(m/2)H2O.............(i)
40ml.........260ml......................0.............0............................(initially)
0................0............................100ml..........................................(finally)
If the condition of pressure and temp. is same, then the ratio of volues of gese would be equal to ratio of their moles(Aogadro law)
Hence the ratio of hydrocrbon and O2=40/260=2/13
but from the equation, the ratio should be =1/(n+m/4)
Hencen+m/4=13/2................................(ii)
Now from the equation we can also see that 1 mole of hydrocarbon would release n mole of CO2 if it completely burns. So the ratio of CO2 and hydrocarbon =n/1=100/40=5/2
hence from (i), we can have value of m and so we can derive the molecular formula for the hydrocarbn.........................